# Decentralized Stochastic Subgradient Methods for Nonsmooth Nonconvex Optimization

### v1

@@ 4.3 Decentralized SGD with gradient-tracking technique
The decentralized SGD method with gradient-tracking technique (DSGD-T) is another widely employed stochastic (sub)gradient method in decentralized learning. Algorithm 3 exhibits the detailed implementations of DSGD-T in minimizing nonsmooth path-differentiable functions. Different from the DSGD in Algorithm 1, DSGD-T introduces a sequence of auxiliary variables to track the average of local subgradients over all the agents. To the best of our knowledge, all the existing works focus on analyzing the convergence of DSGD-T with Clarke regular objective functions, and the convergence of DSGD-T for nonsmooth nonconvex functions remains unexplored. In this subsection, we prove that DSGD-T can be enclosed by our proposed framework (DSM). Therefore, based on our theoretical results exhibited in Theorem 3.8, we can establish the global convergence of DSGD-T in the minimization of nonsmooth path-differentiable functions.
@@
1:Initial point ${\bm{x}}_{0}=[{\bm{x}}_{1,0},\ldots,{\bm{x}}_{d,0}]$, initial descent direction ${\bm{v}}_{0}=[{\bm{v}}_{1,0},\ldots,{\bm{v}}_{d,0}]$ satisfying ${\bm{v}}_{0}{\bm{1}}_{d}=\sum_{i=1}^{d}{\bm{d}}_{0,i}$, where ${\bm{d}}_{i,0}\in\mathcal{D}_{f_{i,j_{i,0}}}({\bm{x}}_{i,0}),$ $j_{i,0}$ is a uniformly random index in $[N_{i}]$, and a mixing matrix ${\bm{W}}$.
@@
6:         Communicate and update local variable ${\bm{x}}_{i,k+1}=\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}-\eta_{k}{\bm{v}}_{i,k}$;
7:         Sample the index $j\in[N_{i}]$ uniformly at random;
8:         Compute ${\bm{d}}_{i,k+1}\in\mathcal{D}_{f_{i,j}}({\bm{x}}_{i,k+1})$;
@@
In the rest of this subsection, we denote ${\bm{V}}_{k}=[{\bm{v}}_{1,k},\ldots,{\bm{v}}_{d,k}]$ at $k$-th iteration in Algorithm 3. Assumption 4.3 and 4.1 are also adopted for DSGD-T. In the following, we condense the assertion that DSGD-T can fit in our framework (DSM), along with convergence analysis, into a single theorem.
@@
Suppose Assumption 4.1 and 4.3 hold for Algorithm 3. Then the iterative formula of $\{{\bm{X}}_{k}\}$ is a specific form of our proposed general framework (DSM). Moreover, for any $i\in[d]$, almost surely, any cluster point of $\{{\bm{x}}_{i,k}:k\in\mathbb{N}\}$ is a $\mathcal{D}_{f}$-stationary point of $f$ and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges.
@@
Let the filtration $\mathcal{F}_{k}=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. By setting ${\bm{Z}}_{k}={\bm{X}}_{k}$, ${\bm{H}}_{k}=\mathbb{E}[{\bm{V}}_{k}|\mathcal{F}_{k}]$ and $\Xi_{k+1}={\bm{V}}_{k}-{\bm{H}}_{k}$, we can conclude that (4.13) coincides with the framework (DSM). Then we aim to check the validity of Assumption 3.1. From the uniform boundedness of $\{{\bm{x}}_{k}\}$ in Assumption 4.3(2), we can conclude that there exists $M_{D}>0$ such that $\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|\leq M_{D}$ almost surely. Then by the update scheme of $\{{\bm{V}}_{k}\}$ in (4.13), it holds that
${\bm{V}}_{k+1}={\bm{D}}_{k+1}{\bm{W}}+\sum_{i=1}^{k}{\bm{D}}_{i}({\bm{W}}-{\bm{I}}_{d}){\bm{W}}^{k+1-i}-{\bm{D}}_{0}{\bm{W}}^{k+1}+{\bm{V}}_{0}{\bm{W}}^{k+1}.$ (4.14)
@@
$\sup_{k\geq 0}\left\|{\bm{V}}_{k}\right\|\leq\sup_{k\geq 0}\left\|{\bm{D}}_{k+1}{\bm{W}}+\sum_{i=1}^{k}{\bm{D}}_{i}({\bm{W}}-{\bm{I}}_{d}){\bm{W}}^{k+1-i}-{\bm{D}}_{0}{\bm{W}}^{k+1}+{\bm{V}}_{0}{\bm{W}}^{k+1}\right\|$ (4.15)
@@
Therefore, we can conclude that $\{{\bm{V}}_{k}\}$ is uniformly bounded almost surely, hence verifying the validity of Assumption 3.1(2). Moreover, the uniform boundedness of $\{\Xi_{k}\}$ and [2, Remark 1.5] illustrates that Assumption 3.1(3) holds almost surely for Algorithm 3. In addition, it is easy to verify that
${\bm{H}}_{k}{\bm{e}}=\mathbb{E}[{\bm{V}}_{k}{\bm{e}}|\mathcal{F}_{k}]=\mathbb{E}[{\bm{D}}_{k}{\bm{e}}|\mathcal{F}_{k}]\in\frac{1}{d}\sum_{i=1}^{d}\mathcal{D}_{f_{i}}({\bm{x}}_{i,k}),\quad\forall k\in\mathbb{N}.$ (4.16)
Therefore, by choosing $\mathcal{H}_{i}=\mathcal{D}_{f_{i}}$ for all $i\in[d]$, we verify the validity of Assumption 3.1(4). Furthermore, according to [8, Section 6] and Remark 4.2, we can assert that $f$ is the Lyapunov function of the differential inclusion
$\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-\mathcal{D}_{f}({\bm{x}}),$ (4.17)
and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$. In addition, the definability of $f$ and $\mathcal{D}_{f}$ illustrates that $\{f({\bm{x}}):{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we verify the validity of Assumption 3.2 with $\mathcal{H}=\mathcal{D}_{f}$, $\Psi=f$, and $\mathcal{A}=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$. Finally, from Theorem 3.8, we can conclude that almost surely, any cluster point of $\{{\bm{x}}_{i,k}:k\in\mathbb{N}\}$ is a $\mathcal{D}_{f}$-stationary point of $f$ and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges. This completes the proof. ∎
@@ 5 Numerical Experiments
In this section, we evaluate the numerical performance of our analyzed decentralized stochastic subgradient methods in Section 4 for decentralized training nonsmooth neural networks. All numerical experiments in this section are conducted on a workstation equipped with two Intel(R) Xeon(R) Processors Gold 5317 CPUs (at 3.00GHz, 18M Cache) and 8 NVIDIA GeForce RTX 4090 GPUs under Ubuntu 20.04.1. We implement all decentralized algorithms with Python 3.8 and PyTorch 1.13.0 using NCCL 2.14.3 (CUDA 11.7) as the communication backend.
5.1 Testing problems and default setting
We implement image classification tasks by the decentralized training of the ResNet-50 [20] network on the CIFAR-10 and CIFAR-100 datasets [25], which both consist of $50000$ training samples and $10000$ test samples. The decentralized network topology is set as ring-structure with $8$ agents, and the mixing matrix is chosen as the Metropolis weight matrix [44]. In fact, in each experiment, we divide the training dataset into 8 equal parts, with each part serving as the local dataset for each agent. During any epoch of training, the local dataset of each agent itself remains unchanged, but the order of batches is reshuffled to ensure compliance with the format of our problem (DOP) and the randomness requirements of all algorithms in Section 4. Moreover, we set the number of epochs as $200$ and choose the batch size as $128$. The strategy for selecting step-size is to decrease it at the $60$-th, $120$-th and $160$-th epochs, with a decaying factor of $0.2$.
5.2 Performance comparison between DSGD, DSGDm, and DSignSGD
We first evaluate the numerical performance of DSGD (Algorithm 1), DSGDm (Algorithm 2 with $\phi=\frac{1}{2}\|{\bm{y}}\|^{2}$), and DSignSGD (Algorithm 2 with $\phi=\left\|{\bm{y}}\right\|_{1}$). All compared methods are executed five times with varying random seeds. In each test instance, all these compared methods are initialized with the same randomly generated initial points. In Algorithm 2, we choose the momentum parameter $\tau$ as $\frac{0.1}{\eta_{0}}$, and the initial descent direction is taken as ${\bm{Y}}_{0}=(1-\tau\eta_{0}){\bm{D}}_{0}{\bm{W}}+\tau\eta_{0}{\bm{D}}_{0}$. Additionally, we set the initial step-size $\eta_{0}$ of DSGD and DSGDm to be $0.2$, and the counterpart of DSignSGD to be $10^{-4}$. We present the numerical results of 3 decentralized subgradient methods in Figure 1, including the test accuracy, train loss, and the consensus error. It is worth mentioning that the consensus error at the $k$-th iteration is measured by $\frac{1}{\sqrt{d}}\|{\bm{X}}_{k}({\bm{I}}_{d}-{\bm{P}})\|_{\mathrm{F}}$. As illustrated in Figure 1, all the compared decentralized stochastic subgradient methods are able to train the ResNet-50 network to a high accuracy. Moreover, we can conclude from Figure 1(a)(d) that DSignSGD achieves similar test accuracy as DSGD and DSGDm, while showing superior performance to DSGD and DSGDm in the aspect of train loss from Figure 1(b)(e). Furthermore, in all the compared methods, the consensus errors diminish to 00 throughout the iterations. These results illustrate that our framework is able to yield efficient stochastic subgradient methods for decentralized training tasks.
(a) Test Accuracy, CIFAR-100
(b) Train Loss, CIFAR-100
(c) Consensus Error, CIFAR-100
(d) Test Accuracy, CIFAR-10
(e) Train Loss, CIFAR-10
(f) Consensus Error, CIFAR-10
@@ 5.2 Performance comparison between DSGD, DSGDm, and DSignSGD
Figure 1: Numerical performance comparison of 3 decentralized subgradient methods in training ResNet50 on CIFAR-10 and CIFAR-100 datasets.
@@ 5.3 Performance comparison between DSGD and DSGD-T
In this subsection, we evaluate the numerical performance of DSGD (Algorithm 1) and DSGD-T (Algorithm 3). Similar to the settings in Section 5.2, we repeat the numerical experiments for $5$ times with varying random seeds for DSGD and DSGD-T. For each random seed, DSGD and DSGD-T are executed with the same initial point ${\bm{x}}_{0}$. The initial step-size is taken as $0.2$ for both algorithms. The numerical results are presented in Figure 2. It can be concluded that DSGD-T demonstrates slightly better performance than DSGD in all aspects, especially in train loss. Therefore, we can conclude that DSGD-T achieves high efficiency in decentralized training tasks while enjoying convergence guarantees from our framework (DSM).
(a) Test Accuracy, CIFAR-100
(b) Train Loss, CIFAR-100
(c) Consensus Error, CIFAR-100
(d) Test Accuracy, CIFAR-10
(e) Train Loss, CIFAR-10
(f) Consensus Error, CIFAR-10
Figure 2: Numerical performance comparison of DSGD and DSGD-T in training ResNet50 on CIFAR-10 and CIFAR-100 datasets.
@@ 6 Conclusion
Decentralized stochastic subgradient methods play an important role in the decentralized training of deep neural networks, where the widely employed nonsmooth building blocks result in non-Clarke regular loss functions. As nearly all of the existing works on decentralized optimization focus on the Clarke regular functions, it is important to provide theoretical guarantees for these decentralized stochastic subgradient methods in the minimization of nonsmooth nonconvex functions. In this paper, we propose a general framework (DSM) to establish the convergence properties of decentralized stochastic subgradient methods for solving (DOP). We show that the sequence of iterates generated by our proposed framework asymptotically approximates the trajectory of its corresponding continuous-time noiseless differential inclusion. Then based on the ODE approaches, we establish the global convergence of our proposed framework (DSM) under mild conditions. We demonstrate the high flexibility of our proposed framework (DSM), in the sense that a wide range of decentralized stochastic subgradient methods fit into our proposed framework (DSM). In particular, we show that the framework (DSM) encloses various popular decentralized stochastic subgradient methods, including DSGD, DSGD-T, DSGDm, and DSignSGD. Based on our theoretical results, we establish global convergence for these decentralized stochastic subgradient methods in the minimization of nonsmooth nonconvex functions, especially in the decentralized training of nonsmooth neural networks. Preliminary numerical experiments verify the validity of our theoretical results and exhibit the high efficiency of these decentralized stochastic subgradient methods enclosed in our framework (DSM). Very recently, apart from the ODE approaches, the closed-measure approaches [4, 9, 26] are developed for establishing the convergence properties of stochastic subgradient methods, with relaxed conditions on the step-sizes and evaluation noises. To the best of our knowledge, all the existing results focus on nonsmooth optimization within a single agent. Hence, how to employ these closed-measure approaches for analyzing decentralized stochastic subgradient methods is an important direction of our future research.

### v2

@@ 4.3 Decentralized SGD with gradient-tracking technique
The decentralized SGD method with gradient-tracking technique (DSGD-T) is another widely employed stochastic (sub)gradient-based method in decentralized learning. Algorithm 3 exhibits the detailed implementations of DSGD-T in minimizing nonsmooth path-differentiable functions. Different from the DSGD in Algorithm 1, DSGD-T introduces a sequence of auxiliary variables to track the average of local subgradients over all the agents. To the best of our knowledge, all the existing works focus on analyzing the convergence of DSGD-T with Clarke regular objective functions, and the convergence of DSGD-T for nonsmooth nonconvex functions remains unexplored.
In this subsection, we prove that DSGD-T can be enclosed by our proposed framework (DSM). Therefore, based on our theoretical results exhibited in Theorem 3.9, we can establish the global convergence of DSGD-T in the minimization of nonsmooth path-differentiable functions.
@@
1:Initial point ${\bm{x}}_{0}=[{\bm{x}}_{1,0},\ldots,{\bm{x}}_{d,0}]$, initial descent direction ${\bm{v}}_{0}=[{\bm{v}}_{1,0},\ldots,{\bm{v}}_{d,0}]$ satisfying ${\bm{v}}_{0}{\bm{1}}_{d}=\sum_{i=1}^{d}{\bm{d}}_{i,0}$, where ${\bm{d}}_{i,0}\in\frac{1}{|\mathcal{B}_{i,0}|}\sum_{l\in\mathcal{B}_{i,0}}D_{F(\cdot;{\bm{s}}_{l})}({\bm{x}}_{i,0}),$ $\mathcal{B}_{i,0}$ is a mini-batch of $\mathcal{S}_{i}$, and a mixing matrix ${\bm{W}}$.
@@
6:         Communicate and update the local variable
${\bm{x}}_{i,k+1}=\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}-\eta_{k}{\bm{v}}_{i,k};$
7:         Randomly select a mini-batch $\mathcal{B}_{i,k+1}\subseteq\mathcal{S}_{i}$;
8:         Compute ${\bm{d}}_{i,k+1}\in\frac{1}{|\mathcal{B}_{i,k+1}|}\sum_{l\in\mathcal{B}_{i,k+1}}D_{F(\cdot;{\bm{s}}_{l})}({\bm{x}})$;
@@
In the rest of this subsection, we denote $,{\bm{V}}_{k}=[{\bm{v}}_{1,k},$$\ldots,{\bm{v}}_{d,k}]$ at $k$-th iteration in Algorithm 3. Assumption 4.3 and 4.1 are also adopted for DSGD-T. In the following, we condense the assertion that DSGD-T can fit in our framework (DSM), along with convergence analysis, into a single theorem.
@@
Suppose Assumption 4.1 and 4.3 hold for Algorithm 3. Then the iterative formula of $\{{\bm{X}}_{k}\}$ is a specific form of our proposed general framework (DSM). Moreover, for any $i\in[d]$, almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $\mathcal{D}_{f}$-stationary point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges.
@@
Let the filtration $\mathcal{F}_{k}=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. By setting ${\bm{Z}}_{k}={\bm{X}}_{k}$, ${\bm{H}}_{k}=\mathbb{E}[{\bm{V}}_{k}|\mathcal{F}_{k}]$ and $\Xi_{k+1}={\bm{V}}_{k}-{\bm{H}}_{k}$, we can conclude that (4.5) coincides with the framework (DSM).
Then we aim to check the validity of Assumption 3.1. From the uniform boundedness of $\{{\bm{X}}_{k}\}$ in Assumption 4.3(1), we can conclude that there exists $M_{D}>0$ such that $\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|\leq M_{D}$ almost surely. Then by the update scheme of $\{{\bm{V}}_{k}\}$ in (4.5), it holds that
${\bm{V}}_{k+1}={}$ ${\bm{D}}_{k+1}{\bm{W}}+\sum_{i=1}^{k}{\bm{D}}_{i}({\bm{W}}-{\bm{I}}_{d}){\bm{W}}^{k+1-i}$
$-{\bm{D}}_{0}{\bm{W}}^{k+1}+{\bm{V}}_{0}{\bm{W}}^{k+1}.$
@@
$\sup_{k\geq 0}\left\|{\bm{V}}_{k}\right\|$
[…]
$-{\bm{D}}_{0}{\bm{W}}^{k+1}+{\bm{V}}_{0}{\bm{W}}^{k+1}\|$
@@
Therefore, we can conclude that $\{{\bm{V}}_{k}\}$ is uniformly bounded almost surely, hence verifying the validity of Assumption 3.1(1).
Moreover, the uniform boundedness of $\{\Xi_{k}\}$ and [30, Remark 1.5] illustrates that Assumption 3.1(2) holds almost surely for Algorithm 3. In addition, followed by $\mathbb{E}[{\bm{d}}_{i,k}|\mathcal{F}_{k}]\in\mathbb{E}_{\mathcal{B}_{i,k}}[D_{\mathcal{B}_{i,k}}({\bm{x}}_{i,k})]=D_{f_{i}}({\bm{x}}_{i,k})$, it is easy to verify that
$\frac{1}{d}{\bm{H}}_{k}{\bm{1}}_{d}$ $=\mathbb{E}[\frac{1}{d}{\bm{V}}_{k}{\bm{1}}_{d}|\mathcal{F}_{k}]=\mathbb{E}[\frac{1}{d}{\bm{D}}_{k}{\bm{1}}_{d}|\mathcal{F}_{k}]$
$\in\frac{1}{d}\sum_{i=1}^{d}\mathcal{D}_{f_{i}}({\bm{x}}_{i,k}),\forall k\in\mathbb{N}.$
Therefore, by choosing $\Phi_{i}=\mathcal{D}_{f_{i}}$ for all $i\in[d]$, we verify the validity of Assumption 3.1(3).
Furthermore, according to [28, Section 6] and Remark 4.2, we can assert that $f$ is the Lyapunov function of the differential inclusion $\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-\mathcal{D}_{f}({\bm{x}})$, and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$. In addition, the definability of $f$ and $\mathcal{D}_{f}$ illustrates that $\{f({\bm{x}}):{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we have verified the validity of Assumption 3.3 with $\Phi=\mathcal{D}_{f}$, $\psi=f$, and $\mathcal{A}=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$.
Finally, from Theorem 3.9, we can conclude that almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $\mathcal{D}_{f}$-stationary point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges. This completes the proof.
@@ 5 Numerical Experiments
(a) Test Accuracy
(b) Train Loss
(c) Consensus Error
@@
Figure 1: Numerical performance comparison of DSGD, DSGD-M and DSignSGD in training ResNet50 on CIFAR-100 dataset.
(a) Test Accuracy
(b) Train Loss
(c) Consensus Error
Figure 2: Numerical performance comparison of DSGD, DSGD-M and DSignSGD in training ResNet50 on CIFAR-10 dataset.
In this section, we evaluate the numerical performance of our analyzed decentralized stochastic subgradient-based methods in Section 4 for decentralized training nonsmooth neural networks. All numerical experiments in this section are conducted on a workstation equipped with two Intel(R) Xeon(R) Processors Gold 5317 CPUs (at 3.00GHz, 18M Cache) and 8 NVIDIA GeForce RTX 4090 GPUs under Ubuntu 20.04.1. We implement all decentralized algorithms with Python 3.8 and PyTorch 1.13.0 using NCCL 2.14.3 (CUDA 11.7) as the communication backend.
5.1 Experiment settings
We conduct a series of image classification tasks with CIFAR-10 and CIFAR-100 datasets [54], which both consist of $50000$ training samples and $10000$ test samples. ResNet-50 [55] is employed as our nonsmooth neural network. The decentralized network topology is set as ring-structure with $8$ agents, and the mixing matrix is chosen as the Metropolis weight matrix [46]. In each experiment, we divide the training dataset into 8 equal parts, with each part serving as the local dataset for each agent. During any epoch of training, the local dataset of each agent itself remains unchanged, but the mini-batch is reshuffled to ensure compliance with the format of our problem (DOP) and the randomness requirements of all algorithms in Section 4.
Moreover, we set the number of epochs as $200$ and choose the batch size as $128$. The strategy for selecting step-size is to decrease it at the $60$-th, $120$-th and $160$-th epochs, with a decaying factor of $0.2$.
5.2 Performance comparison between DSGD, DSGD-M, and DSignSGD
We first evaluate the numerical performance of DSGD (Algorithm 1), DSGD-M (Algorithm 2 with $\phi=\frac{1}{2}\|{\bm{y}}\|^{2}$), and DSignSGD (Algorithm 2 with $\phi=\left\|{\bm{y}}\right\|_{1}$). All compared methods are executed five times with varying random seeds. In each test instance, all these compared methods are initialized with the same randomly generated initial points.
In Algorithm 2, we choose the momentum parameter $\tau$ as $\frac{0.1}{\eta_{0}}$, and the initial descent direction is taken as ${\bm{Y}}_{0}=(1-\tau\eta_{0}){\bm{D}}_{0}{\bm{W}}+\tau\eta_{0}{\bm{D}}_{0}$. Additionally, we set the initial step-size $\eta_{0}$ of DSGD and DSGD-M to be $0.2$, and the counterpart of DSignSGD to be $10^{-4}$.
We present the numerical results of the above-mentioned decentralized stochastic subgradient-based methods in Figure 1 and 2, including the test accuracy, train loss, and consensus error. It is worth mentioning that the consensus error at the $k$-th iteration is measured by $\frac{1}{\sqrt{d}}\|{\bm{X}}_{k}({\bm{I}}_{d}-{\bm{P}})\|_{\mathrm{F}}$. As illustrated in Figure 1 and 2, all the compared decentralized stochastic subgradient-based methods are able to train the ResNet-50 network to a high accuracy. Moreover, we can conclude from Figure 1(a) and 2(a) that DSignSGD achieves similar test accuracy as DSGD and better test accuracy compared to DSGD-M, while showing superior performance to DSGD and DSGD-M in the aspect of train loss from Figure 1(b) and 2(b). Furthermore, in all the compared methods, the consensus errors diminish to 00 throughout the iterations. These results illustrate that our framework is able to yield efficient stochastic subgradient-based methods for decentralized training tasks.
@@ 5.3 Performance comparison between DSGD and DSGD-T
In this subsection, we evaluate the numerical performance of DSGD (Algorithm 1) and DSGD-T (Algorithm 3). Similar to the settings in Section 5.2, we repeat the numerical experiments for $5$ times with varying random seeds for DSGD and DSGD-T. For each random seed, DSGD and DSGD-T are executed with the same initial point ${\bm{x}}_{0}$. The initial step-size is taken as $0.2$ for both algorithms.
The numerical results on the CIFAR-100 and CIFAR-10 datasets are presented in Figure 3 and 4, respectively. It can be concluded that DSGD-T demonstrates slightly better performance than DSGD in all aspects, especially in train loss. Therefore, we can conclude that DSGD-T achieves high efficiency in decentralized training tasks while enjoying convergence guarantees from our framework (DSM).
(a) Test Accuracy
(b) Train Loss
(c) Consensus Error
Figure 3: Numerical performance comparison of DSGD and DSGD-T in training ResNet50 on CIFAR-100 dataset.
(a) Test Accuracy
(b) Train Loss
(c) Consensus Error
Figure 4: Numerical performance comparison of DSGD and DSGD-T in training ResNet50 on CIFAR-10 dataset.
@@ 6 Conclusion
Decentralized stochastic subgradient-based methods play an important role in the decentralized training of deep neural networks, where the widely employed nonsmooth building blocks result in non-Clarke regular loss functions. As nearly all of the existing works on decentralized optimization focus on the Clarke regular functions, there is a gap between theoretical analysis and implementation in real-world training tasks. Additionally, despite the development of various acceleration techniques to improve the performance of decentralized stochastic subgradient methods, the convergence guarantees of these methods in training tasks are largely absent.
To address these issues, we propose a unified framework in (DSM) for developing stochastic subgradient-based methods in nonsmooth decentralized optimization. We conduct consensus analysis and relate the iterates of (DSM) to the trajectories of its corresponding noiseless differential inclusion. Then, under mild conditions, we establish the global convergence of (DSM), in the sense that for any $i,j\in[d]$, $\lim_{k\to+\infty}\left\|z_{i,k}=z_{j,k}\right\|=0$ and $\{z_{i,k}\}$ converges to the stable set of the corresponding noiseless differential inclusion.
Moreover, we show our proposed framework (DSM) is general enough to enclose the implementations of various acceleration techniques for decentralized stochastic subgradient methods, including momentum and gradient-tracking techniques. In particular, we demonstrate that a wide range of decentralized stochastic subgradient-based methods, including DSGD, DSGD-M, and DSGD-T, fits into our proposed framework. Therefore, our theoretical results, for the first time, provide the global convergence guarantees for these decentralized stochastic subgradient-based methods in nonsmooth nonconvex optimization. Furthermore, based on our proposed framework, we develop a new method named DSignSGD, where sign-mapping is employed to regularize the update directions. Preliminary numerical results verify the validity of our developed theoretical results and exhibit the high efficiency of these decentralized stochastic subgradient-based methods enclosed in our framework (DSM).
Our work sets several directions for future research in the field of nonsmooth nonconvex decentralized optimization, including the convergence of decentralized methods with time-varying networks and decentralized methods that incorporate asynchronous communication or compressed communication. Additionally, we are also eager to explore routes that presume relaxed conditions for convergence analysis of decentralized methods in solving (DOP). For instance, closed-measure approaches [51, 52, 35] are developed for establishing the convergence properties of stochastic subgradient-based methods, with relaxed conditions on the step-sizes and evaluation noises. How to employ these closed-measure approaches for analyzing decentralized stochastic subgradient-based methods is worthy of investigating.